# Methods to Characterise Exoplanet Host Stars from Spectroscopy

### v1

@@ 4.1 Fitting observations to synthetic spectra
Computations of a synthetic spectrum requires a model atmosphere based on solutions to the stellar structure equations to synthesize a spectrum. Most stellar atmosphere models are pre-calculated and tabulated on grids describing the profiles of the temperature, surface gravity and abundances as functions of atmospheric depth. Each layer in the model atmosphere is contributing to the formation of absorption line profiles in the final spectrum. Some widely used atmospheric model atmospheres are Atlas12 (Kurucz 2013), Atlas9 (Kurucz 1993a; Heiter et al. 2002), MARCS (Gustafsson et al. 2008) for cool and giant stars, and LL models (Shulyak et al. 2004) for hot main sequence stars. Line lists of atomic and molecular data needed in the computation can be provided by the Vienna Atomic Line Database (VALD3; Piskunov et al. 1995; Ryabchikova et al. 2015).
@@
The dependence of the line profile on the parameter is broadened line wings and a cusp-shaped core. Unfortunately, disentangling the effect on the line profile from and is difficult, leading to a degeneracy between the two. Prior information of may be obtained from time-resolved photometry, available for the known transiting planets, or from asteroseismology (cf. Doyle et al. 2014). If no prior information is available, calibration equations of both turbulent velocities are often used (e.g. Bruntt et al. 2010; Doyle et al. 2014) which allows modelling of . Note that in order to properly model the different velocities above, it is imperative to take into account how the spectrograph itself broadens the lines. Spectral lines can also be pressure broadened through various mechanisms which further broaden the lines, e.g. the Balmer lines. Details of the various broadening mechanisms can be found in many textbooks such as Gray (2008).
@@
There are many softwares that computes synthetic spectra to be used as a model constraint to interpret the observed spectrum. The popular open-source spectroscopic tool iSpec (Blanco-Cuaresma et al. 2014; Blanco-Cuaresma 2019) support several of the most well-known radiative transfer codes such as Spectrocsopy Made Easy (SME; Valenti and Piskunov 1996; Piskunov and Valenti 2017; Wehrhahn et al. 2023), SPECTRUM (Gray and Corbally 1994), Turbospectrum (Plez 2012; de Laverny et al. 2012; Gerber et al. 2023), Synthe/WIDTH9 (Kurucz 1993b), and MOOG (Sneden 1973) where the latter is based on the equivalent width method. The codes often adopt LTE and a plane-parallel geometry as default motivated by the fact that for main-sequence stars, the photosphere constitutes % of the stellar radius. A spherically symmetric geometry is on the other hand required for giant stars where the atmosphere makes up a substantial portion of its radius (Heiter and Eriksson 2006). However, since complex interaction of gas particles and the non-local radiation fields leads to deviations from LTE in the atmospheres of FGKM-type stars, some of the softwares have also the option of non-local thermodynamic equilibrium (NLTE). For instance, SME includes NLTE departure coefficients for the MARCS and LL models atmospheres (Piskunov and Valenti 2017), and Turbospectrum also have the option of NLTE (Gerber et al. 2023).
@@
Fitting can in several of the softwares be made for one or several parameters at the same time using a -minimization algorithm. However, care must be taken when solving for several parameters simultaneously (Torres et al. 2012) due to degeneracies and difficulties in the modelling. Different initial assumptions of one free parameter at a time can therefore be made in the fitting process to iterate to the final solution, thereby mitigating degeneracies. Also, since the surface gravity is often difficult to constrain it is thus advantageous to have external information about the stellar density that can facilitate modelling.
@@
Spectroscopic modelling can take advantage of the sensitivity of certain spectral lines to specific parameters. For instance, all Balmer lines exhibit pressure (collisionally) broadened line wings which makes their profiles strongly sensitive to temperature (e.g. Fuhrmann et al. 1993, 1994; Barklem et al. 2000, 2002) as shown in Fig. 1. The Balmer line wings are formed in the deepest photospheric layers likely close to the LTE. The metal-line blending increase from H at 6562.81 Å to H at 4101.75 Å which perturb the higher transitions of the Balmer line profiles making H the best choice. For late F, G, and early K-stars, H is very insensitive to log and abundance making its line wings an excellent indicator. Two examples of H line profiles towards a K5V and a G0V host star are shown in Fig. 1. It is obvious that the line wings have almost completely disappeared in the K5V spectrum, while prominent and very broad towards the G-star. For M-dwarfs where the line wings of H are absent, TiO lines can also be used as an indicator (Valenti et al. 1998).
@@
For late F- and G-dwarfs there are several pressure-sensitive lines that can be used to constrain log : the Mg I triplet at 5167.33 Å, 5172.70 Å and 5183.62 Å (e.g. Fuhrmann et al. 1997; Valenti and Fischer 2005) as shown in Fig. 2, and the Ca I lines at 6122.23 Å and 6162.18 Å (e.g. Gray 2008). For lower gravity or higher temperature, the line wings disappear due to lower photospheric density or ionisation. The Mg I lines are, however, very wide which makes normalization difficult. Two of the lines lie very close without continuum between them, and there are numerous overlying narrow metal lines in particular for late-type stars. In these cases, only the 5183.62 Å line can be used reliably. Figure 2 shows the Mg I triplet line profiles towards the same host stars as in Fig. 1. As already seen in Fig. 1, the spectral line density is much higher towards the K5V star than the G0V star. However, the effect is even more pronounciated at shorter wavelengths. In addition to surface gravity, the Ca and Mg pressure broadened line wings are, however, to some degree also sensitive to temperature and abundance. The procedure is therefore to model the temperature first, e.g. via H, and the abundance via narrow, unblended Ca (e.g. 6156.02 Å, 6166.439 Å, 6169.042 Å, 6455.985 Å) and Mg (e.g. 5711.09 Å) lines and iterate to a solution.
@@ 4.2 Fitting equivalent widths
In contrast to synthetic spectral synthesis methods, the equivalent width (EW) method begins with the observed spectrum by measuring the strengths of selected absorption lines which are translated into individual line abundances. The method is based on theoretical atmosphere models and excitation and ionisation equilibrium which determines the population in a certain level of an atom or an ion.
The EW of an absorption line is a convenient measurement of its strength. It is defined as the width of a rectangle reaching up to the continuum (with length one) having the same area as the spectral line. Measuring EWs of well-defined weak neutral iron (Fe I) and ionised iron (Fe II) lines is a traditional method to model stellar spectroscopic parameters since there are numerous iron lines in stellar spectra, many with accurate atomic parameters. Metal lines can be very sensitive to temperature although large variations between lines exist. Depending on spectral type, neutral metals are often used as a temperature indicator in solar-type stars, while ionised metal lines are better tracers in early-type stars.
Assuming LTE, the ratio of population in two levels and of an atom (or ion) of a species can be computed with the Bolzmann equation (e.g. Gray 2008; Carroll and Ostlie 2017)
@@
As the temperature increases so will ionisation which occurs quite abruptly once the threshold temperature is reached. In a star’s photosphere, elements exist mainly in just two ionisation stages which can be computed with the Saha equation (e.g. Gray 2008; Carroll and Ostlie 2017) between ionisation states and as
@@
where the electron pressure is (indicator of the surface gravity), is the electron mass, is Planck’s constant, is the partition function, and is the ionisation energy.
@@
The EW method requires measurements of a large number of equivalent widths either measured by direct integration over the entire line or by fitting a Gaussian profile (or a Lorentzian profile that may fit optically thick lines better). Measuring EWs manually is, however, very time-consuming and therefore several softwares have been developed to automate the measurements for example ARES (Sousa et al. 2007) and DAOSPEC (Stetson and Pancino 2008).
@@
Weak and optically thin iron lines depend mainly on and the iron abundance and less on log and . In order to avoid abundance dependence with the EW method, it is best to choose two lines from the same element. However, since continuum normalisation is often a large source of error it may sometimes be necessary to use pairs of lines at nearby wavelengths. In these cases, pairs of similar elements like Fe, V, and Ti that normally have similar abundances, are often used. Different sets of lines are chosen for different spectral types since they are useful in different temperature ranges.
@@
The effective temperature is constrained in the following modelling by the correlation between the excitation potential and the iron abundance of each individual line, the microturbulence is constrained by the correlation between the abundance of each line and the reduced equivalent width, while the surface gravity is constrained by the ionisation balance. The parameters are adjusted until there are no correlations left and all individual abundances are the same (Sousa 2014).
@@
An example of a widely used open radiative transfer code based on the EW method is MOOG (Sneden 1973). This software uses a grid of the ATLAS9 (Kurucz 1993a) plane-parallel model atmospheres to produce model spectra which is compared to the measured EWs of individual Fe I and Fe II lines. The equilibrium conditions are solved simultaneously to derive , log , [Fe/H], and . A good description of the process is found in Sousa (2014). It has been used by for instance Sousa et al. (2021) together with ARES for a homogeneous spectroscopic characterisation of almost one thousand exoplanet host stars (the SWEET-Cat online catalogue). Another example of a software based on the EW method is ODUSSEAS (Observing Dwarfs Using Stellar Spectroscopic Energy-Absorption Shapes; Antoniadis-Karnavas et al. 2020). This code uses the machine learning Python package scikit-learn to offer a quick and automatic derivation of and [Fe/H] for M dwarfs from optical spectroscopy.
@@
When comparing the outcome of the EW and models based on synthetic spectra, and log normally agree within the uncertainties and also with results from asteroseismology within 100 K and 0.1 dex, respectively. The methods, however, do not perform equally well for all spectral types. Since the EW method is based on differential analysis with respect to the Sun, it can be applied to FGK stars with K (Sousa et al. 2011). Modelling based on synthetic spectra and the line shapes are also sensitive to spectral type since some of the fundamental traits, e.g. the broad H line wings, can only be made for solar-type stars since the line wings disappear for colder and hotter stars. In these cases it may still, however, be possible to use other spectral lines (e.g. TiO for low-mass stars). Furthermore, the EW method cannot be used on fast-rotator stars because of the severe line blending. For these stars, it may still be possible to use the synthetic method. Another difference is that the EW method is normally much faster than the synthetic method, while synthetic models provide a more complete description of the star.
@@ 4.3 Empirical methods
Instead of the above methods, it is also possible to compare an observed high-resolution spectrum to spectra of well-characterised stars and in this way derive stellar parameters. An example of such software is Spechmatch-emp (Yee et al. 2017) that compares an observed optical spectrum with a dense empirical spectral library thereby deriving (or log ), , and [Fe/H]. The library contains 404 stars observed with the HIRES instrument with at the Keck telescope. The properties of all the library stars have been derived from asteroseismoloy, interferometry, spectrophotometry and LTE spectral synthesis and represents spectral types from approximately M5 to F1. This method performs very well for the difficult late type stars (K4 stars and later) which are challenging for other methods reaching accuracies of 10 % in stellar radius, 70 K in effective temperature, and 0.09 dex in metallicity ([Fe/H]). The software and the library are publicly available.
@@ 5.1 Radius from spectroscopy and spectral energy distribution
When spectroscopic parameters and photometric measurements are available it is straightforward to obtain the stellar radius from a fit of the observed magnitudes in different bands to its spectral energy distribution (SED). In addition to magnitudes, the SED also depends on the distance (most accurately computed from the observed parallax), the spectroscopic parameters (derived from spectroscopic modelling), the extinction along the line-of-sight, and the stellar radius which is a free parameter in the fit. Parallax measurements have been performed by the European space missions Hipparchos (van Leeuwen 1997) and Gaia (Gaia Collaboration et al. 2016) launched in 1989 and 2013, respectively. The latter mission have provided the community with astrometric and photometric measurements of almost two billion stars in the Milky Way. Gaia has also delivered excellent measurements of the magnitudes in the Gaia optical band. Observations will end in early 2025. Figure 3 shows an example of a SED fit with the Phoenix (Husser et al. 2013) atmospheric model grid. An example of a publicly available software that automatically fits broadband photometry to six different stellar atmosphere models using Nested Sampling algorithms is the ARIADNE (Vines and Jenkins 2022) software.
@@ 5.2 Stellar mass
Unless the star is a binary, the mass cannot usually be determined directly but can be modeled via stellar evolution models. Some of the most popular models are BaSTI (Hidalgo et al. 2018), Padova (Bertelli et al. 2008, 2009), DESP (the Dartmouth Stellar Evolution Database; Dotter et al. 2008), MIST (MESA Isochrones and Stellar Tracks; Choi et al. 2016), PARSEC (the PAdova and TRieste Stellar Evolution Code; Bressan et al. 2012), and Y2 (Yonsei-Yale; Yi et al. 2003; Demarque et al. 2004). A quick way to obtain a Bayesian estimation of both mass and radius based on PARSEC and MESA (Rodrigues et al. 2017) models are the web interfaces Param1.3 and Param1.5 (da Silva et al. 2006). Another publicly available software that provides a simple interface for MCMC fitting of MIST stellar model grids is isochrones (Morton 2015).
A complementing method is to use a set of stars with well-known masses and radii, e.g. through interferometry and eclipsing binaries, to derive empirical calibration equations. Such equations can give the mass and radius of a star given a set of stellar atmosphere values (e.g. Torres et al. 2010; Enoch et al. 2010; Southworth 2011).

### v2

@@ 4.1 Fitting Observations to Synthetic Spectra
Computations of a synthetic spectrum requires a model atmosphere based on solutions to the stellar structure equations to synthesize a spectrum. Most stellar atmosphere models are pre-calculated and tabulated on grids describing the profiles of the temperature, surface gravity, and abundances as functions of atmospheric depth. Each layer in the model atmosphere is contributing to the formation of absorption line profiles in the final spectrum. Some widely used atmospheric model atmospheres are Atlas12 (Kurucz 2013), Atlas9 (Kurucz 1993a; Heiter et al. 2002), MARCS (Gustafsson et al. 2008) for cool and giant stars, and LL models (Shulyak et al. 2004) for hot main sequence stars. Line lists of atomic and molecular data needed in the computation can be provided by the Vienna Atomic Line Database (VALD3; Piskunov et al. 1995; Ryabchikova et al. 2015).
@@
The dependence of the line profile on the parameter is broadened line wings and a cusp-shaped core. Unfortunately, disentangling the effect on the line profile from and is difficult, leading to a degeneracy between the two. Prior information of may be obtained from time-resolved photometry, available for the known transiting planets, or from asteroseismology (cf. Doyle et al. 2014). If no prior information is available, calibration equations of both turbulent velocities are often used (e.g., Bruntt et al. 2010; Doyle et al. 2014) which allows modelling of . Note that in order to properly model the different velocities above, it is imperative to take into account how the spectrograph itself broadens the lines. Spectral lines can also be pressure broadened through various mechanisms which further broaden the lines, e.g., the Balmer lines. Details of the various broadening mechanisms can be found in many textbooks such as Gray (2008).
@@
There are many software that computes synthetic spectra to be used as a model constraint to interpret the observed spectrum. The popular open-source spectroscopic tool iSpec (Blanco-Cuaresma et al. 2014; Blanco-Cuaresma 2019) supports several of the most well-known radiative transfer codes such as Spectroscopy Made Easy (SME; Valenti and Piskunov 1996; Piskunov and Valenti 2017; Wehrhahn et al. 2023), SPECTRUM (Gray and Corbally 1994), Turbospectrum (Plez 2012; de Laverny et al. 2012; Gerber et al. 2023), Synthe/WIDTH9 (Kurucz 1993b), and MOOG (Sneden 1973) where the latter is based on the equivalent width method. The codes often adopt LTE and a plane-parallel geometry as default motivated by the fact that for main sequence stars, the photosphere constitutes % of the stellar radius. A spherically symmetric geometry is on the other hand required for giant stars where the atmosphere makes up a substantial portion of its radius (Heiter and Eriksson 2006). However, since complex interaction of gas particles and the nonlocal radiation fields leads to deviations from LTE in the atmospheres of FGKM-type stars, some of the softwares have also the option of nonlocal thermodynamic equilibrium (NLTE). For instance, SME includes NLTE departure coefficients for the MARCS and LL models atmospheres (Piskunov and Valenti 2017), and Turbospectrum also have the option of NLTE (Gerber et al. 2023).
@@
Fitting can in several of the software be made for one or several parameters at the same time using a -minimization algorithm. However, care must be taken when solving for several parameters simultaneously (Torres et al. 2012) due to degeneracies and difficulties in the modelling. Different initial assumptions of one free parameter at a time can therefore be made in the fitting process to iterate to the final solution, thereby mitigating degeneracies. Also, since the surface gravity is often difficult to constrain, it is thus advantageous to have external information about the stellar density that can facilitate modelling.
@@
Spectroscopic modelling can take advantage of the sensitivity of certain spectral lines to specific parameters. For instance, all Balmer lines exhibit pressure (collisionally) broadened line wings, which makes their profiles strongly sensitive to temperature (e.g., Fuhrmann et al. 1993, 1994; Barklem et al. 2000, 2002) as shown in Fig. 1. The Balmer line wings are formed in the deepest photospheric layers likely close to the LTE. The metal-line blending increases from H at 6562.81 Å to H at 4101.75 Å, which perturbs the higher transitions of the Balmer line profiles making H the best choice. For late F, G, and early K-stars, H is very insensitive to log and abundance making its line wings an excellent indicator. Two examples of H line profiles toward a K5V and a G0V host star are shown in Fig. 1. It is obvious that the line wings have almost completely disappeared in the K5V spectrum, while prominent and very broad toward the G-star. For M-dwarfs where the line wings of H are absent, TiO lines can also be used as a indicator (Valenti et al. 1998).
@@
For late F- and G-dwarfs there are several pressure-sensitive lines that can be used to constrain log : The Mg I triplet at 5167.33 Å, 5172.70 Å, and 5183.62 Å (e.g., Fuhrmann et al. 1997; Valenti and Fischer 2005) as shown in Fig. 2, and the Ca I lines at 6122.23 Å and 6162.18 Å (e.g. Gray 2008). For lower gravity or higher temperature, the line wings disappear due to lower photospheric density or ionization. The Mg I lines are, however, very wide which makes normalization difficult. Two of the lines lie very close without continuum between them, and there are numerous overlying narrow metal lines in particular for late-type stars. In these cases, only the 5183.62 Å line can be used reliably. Figure 2 shows the Mg I triplet line profiles toward the same host stars as in Fig. 1. As already seen in Fig. 1, the spectral line density is much higher toward the K5V star than the G0V star. However, the effect is even more pronounciated at shorter wavelengths. In addition to surface gravity, the Ca and Mg pressure broadened line wings are, however, to some degree also sensitive to temperature and abundance. The procedure is therefore to model the temperature first, e.g., via H, and the abundance via narrow, unblended Ca (e.g., 6156.02 Å, 6166.439 Å, 6169.042 Å, 6455.985 Å) and Mg (e.g., 5711.09 Å) lines and iterate to a solution.
@@ 4.2 Fitting Equivalent Widths
In contrast to synthetic spectral synthesis methods, the equivalent width (EW) method begins with the observed spectrum by measuring the strengths of selected absorption lines which are translated into individual line abundances. The method is based on theoretical atmosphere models and excitation and ionization equilibrium which determines the population in a certain level of an atom or an ion.
The EW of an absorption line is a convenient measurement of its strength. It is defined as the width of a rectangle reaching up to the continuum (with length one) having the same area as the spectral line. Measuring EWs of well-defined weak neutral iron (Fe I) and ionized iron (Fe II) lines is a traditional method to model stellar spectroscopic parameters since there are numerous iron lines in stellar spectra, many with accurate atomic parameters. Metal lines can be very sensitive to temperature although large variations between lines exist. Depending on spectral type, neutral metals are often used as a temperature indicator in solar-type stars, while ionized metal lines are better tracers in early-type stars.
Assuming LTE, the ratio of population in two levels and of an atom (or ion) of a species can be computed with the Boltzmann equation (e.g., Gray 2008; Carroll and Ostlie 2017)
@@
As the temperature increases so will ionization which occurs quite abruptly once the threshold temperature is reached. In a star’s photosphere, elements exist mainly in just two ionization stages, which can be computed with the Saha equation (e.g., Gray 2008; Carroll and Ostlie 2017) between ionization states and as
@@
where the electron pressure is (indicator of the surface gravity), is the electron mass, is Planck’s constant, is the partition function, and is the ionization energy.
@@
The EW method requires measurements of a large number of equivalent widths either measured by direct integration over the entire line or by fitting a Gaussian profile (or a Lorentzian profile that may fit optically thick lines better). Measuring EWs manually is, however, very time-consuming and therefore several software have been developed to automate the measurements for example ARES (Sousa et al. 2007) and DAOSPEC (Stetson and Pancino 2008).
@@
Weak and optically thin iron lines depend mainly on and the iron abundance and less on log and . In order to avoid abundance dependence with the EW method, it is best to choose two lines from the same element. However, since continuum normalization is often a large source of error it may sometimes be necessary to use pairs of lines at nearby wavelengths. In these cases, pairs of similar elements like Fe, V, and Ti that normally have similar abundances, are often used. Different sets of lines are chosen for different spectral types since they are useful in different temperature ranges.
@@
The effective temperature is constrained in the following modelling by the correlation between the excitation potential and the iron abundance of each individual line, the microturbulence is constrained by the correlation between the abundance of each line and the reduced equivalent width, while the surface gravity is constrained by the ionization balance. The parameters are adjusted until there are no correlations left and all individual abundances are the same (Sousa 2014).
@@
An example of a widely used open radiative transfer code based on the EW method is MOOG (Sneden 1973). This software uses a grid of the ATLAS9 (Kurucz 1993a) plane-parallel model atmospheres to produce model spectra which is compared to the measured EWs of individual Fe I and Fe II lines. The equilibrium conditions are solved simultaneously to derive , log , [Fe/H], and . A good description of the process is found in Sousa (2014). It has been used by for instance Sousa et al. (2021) together with ARES for a homogeneous spectroscopic characterization of almost one thousand exoplanet host stars (the SWEET-Cat online catalogue). Another example of a software based on the EW method is ODUSSEAS (Observing Dwarfs Using Stellar Spectroscopic Energy-Absorption Shapes; Antoniadis-Karnavas et al. 2020). This code uses the machine learning Python package scikit-learn to offer a quick and automatic derivation of and [Fe/H] for M dwarfs from optical spectroscopy.
@@
When comparing the outcome of the EW and models based on synthetic spectra, and log normally agree within the uncertainties and also with results from asteroseismology within 100 K and 0.1 dex, respectively. The methods, however, do not perform equally well for all spectral types. Since the EW method is based on differential analysis with respect to the Sun, it can be applied to FGK stars with K (Sousa et al. 2011). Modelling based on synthetic spectra and the line shapes is also sensitive to spectral type since some of the fundamental traits, e.g., the broad H line wings, can only be made for solar-type stars since the line wings disappear for colder and hotter stars. In these cases it may still, however, be possible to use other spectral lines (e.g. TiO for low-mass stars). Furthermore, the EW method cannot be used on fast-rotator stars because of the severe line blending. For these stars, it may still be possible to use the synthetic method. Another difference is that the EW method is normally much faster than the synthetic method, while synthetic models provide a more complete description of the star.
@@ 4.3 Empirical Methods
Instead of the above methods, it is also possible to compare an observed high-resolution spectrum to a spectra of well-characterized stars and in this way derive stellar parameters. An example of such software is Spechmatch-emp (Yee et al. 2017) that compares an observed optical spectrum with a dense empirical spectral library thereby deriving (or log ), , and [Fe/H]. The library contains 404 stars observed with the HIRES instrument with at the Keck telescope. The properties of all the library stars have been derived from asteroseismology, interferometry, spectrophotometry, and LTE spectral synthesis and represent spectral types from approximately M5 to F1. This method performs very well for the difficult late type stars (K4 stars and later) which are challenging for other methods reaching accuracies of 10 % in stellar radius, 70 K in effective temperature, and 0.09 dex in metallicity ([Fe/H]). The software and the library are publicly available.
@@ 5.1 Radius from Spectroscopy and Spectral Energy Distribution
When spectroscopic parameters and photometric measurements are available it is straightforward to obtain the stellar radius from a fit of the observed magnitudes in different bands to its spectral energy distribution (SED). In addition to magnitudes, the SED also depends on the distance (most accurately computed from the observed parallax), the spectroscopic parameters (derived from spectroscopic modelling), the extinction along the line of sight, and the stellar radius which is a free parameter in the fit. Parallax measurements have been performed by the European space missions Hipparchos (van Leeuwen 1997) and Gaia (Gaia Collaboration et al. 2016) launched in 1989 and 2013, respectively. The latter mission has provided the community with astrometric and photometric measurements of almost two billion stars in the Milky Way. Gaia has also delivered excellent measurements of the magnitudes in the Gaia optical band. Observations will end in early 2025. Figure 3 shows an example of a SED fit with the Phoenix (Husser et al. 2013) atmospheric model grid. An example of a publicly available software that automatically fits broadband photometry to six different stellar atmosphere models using Nested Sampling algorithms is the ARIADNE (Vines and Jenkins 2022) software.
@@ 5.2 Stellar Mass
Unless the star is a binary, the mass cannot usually be determined directly but can be modelled via stellar evolution models. Some of the most popular models are BaSTI (Hidalgo et al. 2018), Padova (Bertelli et al. 2008, 2009), DESP (the Dartmouth Stellar Evolution Database; Dotter et al. 2008), MIST (MESA Isochrones and Stellar Tracks; Choi et al. 2016), PARSEC (the PAdova and TRieste Stellar Evolution Code; Bressan et al. 2012), and Y2 (Yonsei-Yale; Yi et al. 2003; Demarque et al. 2004). A quick way to obtain a Bayesian estimation of both mass and radius based on PARSEC and MESA (Rodrigues et al. 2017) models are the web interfaces Param1.3 and Param1.5 (da Silva et al. 2006). Another publicly available software that provides a simple interface for MCMC fitting of MIST stellar model grids is isochrones (Morton 2015).
A complementing method is to use a set of stars with well-known masses and radii, e.g. through interferometry and eclipsing binaries, to derive empirical calibration equations. Such equations can give the mass and radius of a star given a set of stellar atmosphere values (e.g., Torres et al. 2010; Enoch et al. 2010; Southworth 2011).